OpenStudy (anonymous):

Evaluate m/n times p for M = 10, N = -5, and P = -2 .

OpenStudy (anonymous):

10 divided by -5 is?

Mehek (mehek14):

\(\dfrac{10}{-5}*-2\)

OpenStudy (anonymous):

-2

Mehek (mehek14):

divide \(\dfrac{10}{-5}\)and then multiply what you get by \(-2\)

Mehek (mehek14):

ok you got -2 what is -2*-2

OpenStudy (anonymous):

and -2 times -2 is? thats your answer :)

OpenStudy (anonymous):

4 is not a option @.Gjallarhorn.

Mehek (mehek14):

what're the options?

OpenStudy (anonymous):

thats odd

OpenStudy (anonymous):

-100 100 -1 1

OpenStudy (anonymous):

woah what

Mehek (mehek14):

0_0

Mehek (mehek14):

you sure you're looking at the right question?

OpenStudy (anonymous):

yea i had to type out the m/n time p part... its got the m then a line under the m with n with a centered dot p

OpenStudy (anonymous):

oh i got the answer then

OpenStudy (anonymous):

so m is on top and n times p on the bottom

Mehek (mehek14):

\(\dfrac{10}{-5*-2}\)

Mehek (mehek14):

you would multiply -5*-2

Mehek (mehek14):

and then divide by 10

OpenStudy (anonymous):

yea like that

Mehek (mehek14):

so multiply the 2 numbers in the denominator and then divide

OpenStudy (anonymous):

answer is 1 then

OpenStudy (anonymous):

so its 1

OpenStudy (anonymous):

okay thanks @Mehek14 @.Gjallarhorn.